# A Communication Consistent Approach to Signal Temporal Logic Task Decomposition in Multi-Agent Systems

### v2

@@ Abstract
We address the problem of decomposing a global task assigned to a multi-agent system, where the task is expressed using a fragment of Signal Temporal Logic (STL) and communication among agents is range-limited. The global task is expressed as the conjunction of local tasks over individual and relative agent’s states, thus naturally inducing the definition of a task graph. Due to the limited communication range, inconsistencies among the edges of the communication and task graph prevent the application of previously derived state feedback control laws for the satisfaction of global STL tasks. To resolve this issue, we propose a task decomposition mechanism reassigning tasks between non-communicating agents to communicating ones, thus ensuring consistency between task and communication graphs. By assuming that the STL tasks are defined over concave predicate functions with polytopic super-level sets, the decomposition can be framed as a parameter optimization problem, solvable via decentralized convex optimization. To guarantee the soundness of our approach, we present various conditions under which the tasks defined in the applied STL fragment are unsatisfiable, and we derive sufficient conditions such that our decomposition approach yields satisfiable global tasks after decomposition.
{IEEEkeywords}
Signal Temporal Logic, Multi-Agent Systems, Optimization.
1 Introduction
@@
Recently, Signal Temporal Logic (STL) has received increasing attention as a tool to monitor and control the run-time behaviour of multi-agent cyber-physical systems [1]. Notably, an extensive body of literature has been developed dealing with synthesizing admissible plans and feedback solutions that achieve the satisfaction of a given global STL task assigned to a multi-agent system (MAS) [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. While the scalability of the proposed approaches varies depending on the semantic richness of the considered task and the number of agents involved in the task, task decomposition has been recently explored as a solution to mitigate the curse of dimensionality at the control and planning level, borrowing from similar approaches applied for other types of temporal logics [12, 13]. Namely, the authors in [14, 15, 16, 17] first proposed optimization-based approaches to split complex global STL tasks into simpler ones assigned to teams of agents in the systems, whose satisfaction, when jointly considered, recovers the satisfaction of the original global task. However, the influence of agent-to-agent communication has received only minor attention in these settings [18, 19, 20, 21]. If, loosely speaking, we denote task dependencies to be consistent when the agents involved in the task are in communication with each other, then such consistency is a key enabler for many of the previous control/planning approaches. From this observation, the contribution of this work is to establish an approach whereby agents can, in a decentralized fashion, reassign task dependencies that are not consistent with the range-limited communication constraints while recovering the original task satisfaction.
@@
We consider the setting in which the global STL task is expressed as a conjunction of local tasks assigned to single agents and pairs of agents in the system. These kinds of specifications are particularly suited for time-varying formation control of multi-agent systems, which find wide applications in robotics in the context of surveillance and reconnaissance tasks, localization and exploration tasks, and coverage control, to mention a few [22, 23, 24, 25]. Within this framework, task dependencies among pairs of agents are captured by a task graph. At the same time, range-limited communication naturally induces the definition of a communication graph where edges are defined by pairs of communicating agents. We propose to decompose tasks assigned over pairs of non-communicating agents as a conjunction of tasks defined over agents forming a multi-hop communication path between the former non-communicating pair. More specifically, we show that by considering concave predicate functions whose super-level set is defined by a bounded polytope, our decomposition mechanism can be cast as a convex optimization problem. In addition, we show that such optimization can be efficiently decentralized over the edges of the communication graph when its structure is acyclic, thus providing scalability of the proposed approach. Eventually, we provide sufficient conditions under which the new task derived from the decomposition is satisfiable and its satisfaction implies the original global task. To the best of our knowledge, this work represents the first decentralized solution to an STL task decomposition problem in multi-agent settings.
@@
Previous literature in STL task decomposition includes the work by [17], which first introduces a continuous-time and continuous-space STL task decomposition via convex optimization in a centralized fashion. On the other hand, the authors in [14, 15] employ a Mixed-Integer Linear Program (MILP) formulation to decompose a global task expressed in a variant of the STL formalism, denoted as Capability STL (CaSTL), as local team-wise tasks. Differently from [17, 14, 15], our decomposition approach considers communication constraints at the task decomposition level. It is noted however, that the STL fragment considered in [14, 15] is syntactically richer than the one considered in [17] and our work. The relevance of this work should be considered with respect to previous literature in feedback control solutions for the satisfaction of STL tasks in multi-agent settings, where all-to-all agent communication is explicitly or implicitly assumed [26, 27, 28, 19]. Namely, we aim to mitigate the complete connectivity requirement (i.e. every agent is directly connected to every other agent) of these previously developed feedback control solutions, while also retaining their adaptive and reactive properties, by decomposing tasks defined over non-communicating agents into conjunctions of tasks over those that do communicate. We list the following contributions:
@@
iii) We provide a set of conditions under which conjunctions of collaborative tasks over the proposed task graph are unsatisfiable (referenced as conflicting conjunctions hereafter).
@@
Compared to our work in [21], the novel contribution of this work is to show that the task decomposition can be achieved in a decentralized manner and over a broader class of predicate types. The paper is organized as follows: Section 2 reviews polytope operations, STL, and graph theory. The problem statement is provided in Section 3. Section 4 introduces the concept of conflicting conjunctions of collaborative tasks, while Sections 5-6 develop the task decomposition approach. Simulations and conclusions are presented in Sections 7 and 8.
2 Notation and Preliminaries
@@
Let $\mathcal{V}=\{1,\ldots N\}$ be the indices of the agents in a multi-agent system, each following the nonlinear dynamics:
$\dot{\bm{x}}_{i}=f_{i}(\bm{x}_{i},\bm{u}_{i}),\;$ (1)
where $\bm{x}_{i}\in\mathbb{X}_{i}\subset\mathbb{R}^{n}$ and $\bm{u}_{i}\in\mathbb{U}_{i}\subset\mathbb{R}^{m_{i}}$ represent the state and control input for agent $i$ with $m_{i}\geq 1$. Without loss of generality, consider $\mathbb{X}_{i}$ and $\mathbb{U}_{i}$ to be compact sets containing the origin and let $f_{i}:\mathbb{X}_{i}\times\mathbb{U}_{i}\rightarrow\mathbb{R}^{n}$ be locally Lipschitz continuous on $\mathbb{X}_{i}$ and $\mathbb{U}_{i}$. Furthermore, let $\bm{u}_{i}:\mathbb{R}_{+}\rightarrow\mathbb{U}_{i}\in\mathcal{U}_{i}$ and $\bm{x}_{i}:\mathbb{R}_{+}\rightarrow\mathbb{X}_{i}\in\mathcal{X}_{i}$ represent the input and state signals for each agent $i$ where $\mathcal{U}_{i}$ is the set of Lipschitz continuous input signals and $\mathcal{X}_{i}$ is the set of absolutely continuous solutions of (1) under the input signals from $\mathcal{U}_{i}$. The complete state of the MAS is then given by $\bm{x}:=[\bm{x}_{i}]_{i\in\mathcal{V}}\in\mathbb{X}$, with $\mathbb{X}:=\bigtimes_{i\in\mathcal{V}}\mathbb{X}_{i}$, and similarly $\bm{x}(t)=[\bm{x}_{i}(t)]_{i\in\mathcal{V}}\in\mathcal{X}$ where $\mathcal{X}:=\bigtimes_{i\in\mathcal{V}}\mathcal{X}_{i}$. Let $\bm{e}_{ij}:=\bm{x}_{j}-\bm{x}_{i}\in\mathbb{X}_{ij}$ represent the relative state vector for each $i,j\in\mathcal{V}$ where $\mathbb{X}_{ij}:=\mathbb{X}_{j}\ominus\mathbb{X}_{i}$ and let $S\in\mathbb{R}^{n_{p}\times n}$ be a selection matrix [^1] A selection matrix applies to select $m\leq n$ unique elements from a vector of dimensions $n$ such that $S\in\mathbb{R}^{m\times n}$ with $S[i,j]\in\{0,1\}$; $\sum_{j=1}^{n}S[i,j]=1,\,\forall i=1,\ldots m$ and $\sum_{i=1}^{m}S[i,j]=1,\,\forall j=1,\ldots n$. such that $\bm{p}_{i}=S\bm{x}_{i},\;\forall i\in\mathcal{V}$, represents the position vector of agent $i$ with $n_{p}\leq n$.
2.1 Polytopes and their properties
In this subsection, some fundamental properties of polytopes based on [32, Ch. 0-1] and [33, Sec. 3] are revised and adapted to the settings of this work for clarity of presentation. Results not directly available from these references are proven in [34].
@@ 2.1 Polytopes and their properties
The inequality in the definition of $\mathcal{P}(A,\bm{c},\bm{z})$ should be interpreted row-wise. Moreover, the centre vector $\bm{c}$ is introduced for the convenience of analysis, but the polytope definition $\mathcal{P}(A,\tilde{z})=\{\bm{x}\in\mathbb{R}^{n}|A\bm{x}\leq\tilde{\bm{z}}\}$ in [32, pp. 28] is equivalent to ours letting $\tilde{\bm{z}}=\bm{z}+A\bm{c}$.
@@
( [32, pp. 4]) For a finite set of points $V=\{\bm{v}_{k}\}_{k=1}^{|V|}\subset\mathbb{R}^{n}$ with $|V|\geq 1$, the convex hull of $V$ is defined as $conv(V):=\{\sum_{k=1}^{|V|}\lambda_{k}\bm{v}_{k}|\sum_{k=1}^{|V|}\lambda_{k}=1,\,\lambda_{k}\geq 0,\,\forall k=1,\ldots|V|\}$.
@@
Proposition 1 establishes three equivalent representations for a polytope $\mathcal{P}(A,\bm{c},\bm{z})$. Given $A$, $\bm{z}$ and $\bm{c}$, then a generator set $V$ can be obtained algorithmically (see for example [35]), while the functional representation $h(\bm{x})$ is obtained at no additional computational cost [32]. While for a given polytope the generator set $V$ is not unique, we define $\nu:\mathcal{P}(A,\bm{c},\bm{z})\mapsto V^{*}$ as the map between a polytope $\mathcal{P}(A,\bm{c},\bm{z})$ and the set of unique generators with minimal cardinality $V^{*}$ (the set of vertices for $\mathcal{P}$ [32, Proposition 2.2]). We next define when two polytopes are said to be similar and what properties can be derived from this similarity relation.
@@
( [34]) Let the polytope $\mathcal{P}(\bm{A},0,\bm{z})$, the center vector $\bm{c}\in\mathbb{R}^{n}$ and scale $\alpha>0$ such that $\bm{\eta}=[\bm{c},\alpha]\in\mathbb{R}^{n+1}$. Moreover, let $V=\nu(\mathcal{P}(A,\bm{0},\bm{z}))$ be the generator set, such that $V=\{\bm{v}_{k}\}_{k=1}^{|V|}$, then
@@
In other words, the generators of a polytope $\mathcal{P}(A,\bm{c},\alpha\bm{z})$ are directly computed by translating and scaling the generators of the similar polytope $\mathcal{P}(A,\bm{0},\bm{z})$ (see Fig. 1). The next proposition extends the result of Proposition 2 to the Minkowski sum of similar polytopes.
@@
( [34]) Let $\mathcal{P}(A,\bm{c}_{i},\alpha_{i}\bm{z}),\,\forall i\in\mathcal{I}$ be similar polytopes for some index set $\mathcal{I}$ and $\alpha_{i}\in\mathbb{R}_{+}$, then
@@
Next, Proposition 4 provides a set of linear inequalities applicable to check the intersection and inclusion of polytopes.
@@
( [34]) Let the polytopes $\mathcal{P}(A_{1},\bm{c}_{1},\alpha_{1}\bm{z}_{1})$, $\mathcal{P}(A_{2},\bm{c}_{2},\alpha_{2}\bm{z}_{2})$ such that $A_{1}\in\mathbb{R}^{m_{1}\times n},A_{2}\in\mathbb{R}^{m_{2}\times n}$ and let $\bm{\eta}_{1}=[\bm{c}_{1},\alpha_{1}]$ and $\bm{\eta}_{2}=[\bm{c}_{2},\alpha_{2}]$. Furthermore, let $V_{1}=\nu(\mathcal{P}(A_{1},\bm{0},\bm{z}_{1})=\{\bm{v}_{k}\}_{k=1}^{|V_{1}|}$. Then, the inclusion relation $\mathcal{P}(A_{1},\bm{c}_{1},\alpha_{1}\bm{z}_{1})\subseteq\mathcal{P}(A_{2},\bm{c}_{2},\alpha_{2}\bm{z}_{2})$ holds if and only if
@@
The section is concluded with the following result, whose relevance is clarified in Sec. 5.1.
@@
Thus Proposition 5 supports the intuition that given a polytope and a Minkowski sum of polytopes similar to the former, then the Minkowski sum is included in the original polytope only if the sum of the scale factors $\alpha_{i}$ is less than or equal to 1. The relevance of this fact is clarified in Sec. 5.1.
2.2 Signal Temporal Logic
Signal Temporal Logic (STL) is a predicate logic applied to formally define spatio-temporal behaviours (tasks) for state signals $\bm{x}(t)\in\mathcal{X}$. Let $h:\mathcal{D}\rightarrow\mathbb{R}$, with $\mathcal{D}\subseteq\mathbb{X}$ be a predicate function and let the boolean-valued predicate $\mu(h(\bm{x})):=\left\{\begin{smallmatrix}\top&\text{if}\;h(\bm{x})\geq 0\\ \bot&\text{if}\;h(\bm{x})<0\end{smallmatrix}\right$.. Then, the recursively defined STL grammar applies to construct complex specifications from a set of elementary predicates as

### v3

@@ Abstract
We consider the problem of decomposing a global task assigned to a multi-agent system, expressed as a formula within a fragment of Signal Temporal Logic (STL), under range-limited communication. Given a global task expressed as a conjunction of local tasks defined over the individual and relative states of agents in the system, we propose representing task dependencies among agents as edges of a suitably defined task graph. At the same time, range-limited communication naturally induces the definition of a communication graph that defines which agents have access to each other’s states. Within these settings, inconsistencies arise when a task dependency between a pair of agents is not supported by a corresponding communication link due to the limited communication range. As a result, state feedback control laws previously derived to achieve the tasks’ satisfaction can not be leveraged. We propose a task decomposition mechanism to distribute tasks assigned to pairs of non-communicating agents in the system as conjunctions of tasks defined over the relative states of communicating agents, thus enforcing consistency between task and communication graphs. Assuming the super-level sets of the predicate functions composing the STL tasks are bounded polytopes, our task decomposition mechanism can be cast as a parameter optimization problem and solved via state-of-the-art decentralized convex optimization algorithms. To guarantee the soundness of our approach, we present various conditions under which the tasks defined in the applied STL fragment are unsatisfiable, and we show sufficient conditions such that our decomposition approach yields satisfiable global tasks after decomposition.
I Introduction
@@
Recently, Signal Temporal Logic (STL) has received increasing attention as a tool to monitor and control the run-time behaviour of multi-agent cyber-physical systems [1]. Notably, an extensive body of literature has been developed dealing with synthesizing admissible plans and feedback solutions that achieve the satisfaction of a given global STL task assigned to a multi-agent system [2, 3, 4, 5, 6, 7, 8, 9, 10, 11]. While the scalability of the proposed approaches varies depending on the semantic richness of the considered task and the number of agents involved in the task, task decomposition has been recently explored as a solution to mitigate the curse of dimensionality at the control and planning level, borrowing from similar approaches applied for other types of temporal logics [12, 13]. Namely, the authors in [14, 15, 16, 17] first proposed optimization-based approaches to split complex global STL tasks into simpler ones assigned to teams of agents in the systems, whose satisfaction, when jointly considered, recovers the satisfaction of the original global task. However, the influence of range-limited communication has received only minor attention in these settings [18, 19, 20, 21]. Indeed, in environments where all-to-all communication can not be ensured, consistency between task dependencies and communication is the key factor enabling many of the previous control/planning approaches to be reactive to changes in the environment. From this observation, the contribution of this work is to establish an approach whereby agents can, in a decentralized fashion, reassign task dependencies that are not consistent with the range-limited communication constraints while recovering the original task satisfaction.
@@
We consider the setting in which the global STL task is expressed as a conjunction of local tasks assigned to single agents and pairs of agents in the system. These kinds of specifications are particularly suited for time-varying formation control of multi-agent systems, which find wide applications in robotics in the context of surveillance and reconnaissance tasks, localization and exploration tasks, and coverage control, to mention a few [22, 23, 24, 25]. Within this framework, task dependencies among pairs of agents are captured by a task graph. At the same time, range-limited communication naturally induces the definition of a communication graph where edges are defined by pairs of communicating agents. We propose to decompose tasks assigned over pairs of non-communicating agents as a conjunction of tasks defined over agents forming a multi-hop communication path between the former. More specifically, we show that by considering concave predicate functions whose super-level set is defined by a bounded polytope, our decomposition mechanism can be cast as a convex optimization problem. In addition, we show that such optimization can be efficiently decentralized over the edges of the communication graph when its structure is forced to be acyclic by letting each agent select a limited number of communication links, thus providing scalability of the proposed approach. Eventually, we provide sufficient conditions under which the new task derived from the decomposition is satisfiable and its satisfaction implies the original global task. To the best of our knowledge, this work represents the first decentralized solution to an STL task decomposition problem in multi-agent settings.
@@
Previous literature in STL task decomposition includes the work by [17], which first introduces a continuous-time and continuous-space STL task decomposition via convex optimization, which we build upon. On the other hand, the authors in [14, 15] employ a Mixed-integer Linear Program (MILP)s formulation to decompose tasks expressed in a variant of the STL formalism, denoted as Capability STL (CaSTL), as local team-wise tasks. Differently from these previous works, our decomposition approach considers communication constraints at the task decomposition level. It is noted however that the STL fragment considered in [14, 15] is syntactically richer than the one considered in [17] and our work. The relevance of this work should be considered with respect to previous reactive feedback control solutions for the satisfaction of STL tasks in multi-agent settings, where all-to-all agent communication is explicitly or implicitly assumed [26, 27, 28, 19]. Namely, we aim to mitigate the full connectivity requirement for this previous literature by decomposing tasks defined over non-communicating agents into conjunctions of tasks over those that do communicate through the communication network. Since decomposition occurs offline, our method retains the reactive, adaptive qualities of previous feedback control laws in range-limited communication settings. We list the following contributions:
@@
iii) We provide a set of conditions under which conjunctions of collaborative tasks over the proposed task graph are unsatisfiable (conflicting conjunctions hereafter).
@@
The rest of the paper is organized as follows: Section II reviews polytope operations, STL, and graph theory. The problem statement is provided in Section III. Section IV introduces conflicting conjunctions in collaborative tasks, while Sections V-VI develop the task decomposition approach. Numerical simulations and conclusions are presented in Sections VII and VIII.
II Notation and Preliminaries
@@
Let $\mathcal{V}=\{1,\ldots N\}$ be the set of indices assigned to each agent in a multi-agent system and let each agent be governed by the input-affine nonlinear dynamics:
$\dot{\bm{x}}_{i}=f_{i}(\bm{x}_{i})+g_{i}(\bm{x}_{i})\bm{u}_{i},\;$ (1)
where $\bm{x}_{i}\in\mathbb{X}_{i}\subset\mathbb{R}^{n}$ and $\bm{u}_{i}\in\mathbb{U}_{i}\subset\mathbb{R}^{m_{i}}$ represent the state and control input for agent $i$ with input dimensions $m_{i}\geq 1$. Without loss of generality, consider $\mathbb{X}_{i}$ and $\mathbb{U}_{i}$ to be compact sets containing the origin. Let $f_{i}:\mathbb{X}_{i}\rightarrow\mathbb{R}^{n}$, $g_{i}:\mathbb{X}_{i}\rightarrow\mathbb{R}^{n\times m_{i}}$ be locally Lipschitz continuous functions on $\mathbb{X}_{i}$. Furthermore, let $\bm{u}_{i}:\mathbb{R}_{+}\rightarrow\mathbb{U}_{i}\in\mathcal{U}_{i}$ and $\bm{x}_{i}:\mathbb{R}_{+}\rightarrow\mathbb{X}_{i}\in\mathcal{X}_{i}$ represent the input and state signals for each agent $i$ such that $\mathcal{U}_{i}$ is the set of Lipschitz continuous input signals and $\mathcal{X}_{i}$ is the set of absolutely continuous solutions of (1) under the input signals from $\mathcal{U}_{i}$. The global MAS dynamics is then compactly written as
$\dot{\bm{x}}=f(\bm{x})+g(\bm{x})\bm{u},$ (2)
with $\bm{x}:=[\bm{x}_{i}]_{i\in\mathcal{V}}\in\mathbb{X},\bm{u}:=[\bm{u}_{i}]_{i\in\mathcal{V}}\in\mathbb{U},f(\bm{x}):=[f_{i}(\bm{x}_{i})]_{i\in\mathcal{V}}$, $g(\bm{x}):=blk(g_{1}(\bm{x}_{1}),\ldots g_{N}(\bm{x}_{N}))$, $\mathbb{X}:=\bigtimes_{i\in\mathcal{V}}\mathbb{X}_{i}$ and $\mathbb{U}:=\bigtimes_{i\in\mathcal{V}}\mathbb{U}_{i}$. Consistently, the definition of state and input signals for the MAS are given as $\bm{x}(t)=[\bm{x}_{i}(t)]_{i\in\mathcal{V}}\in\mathcal{X}$ and $\bm{u}:=[\bm{u}_{i}(t)]_{i\in\mathcal{V}}\in\mathcal{U}$ with $\mathcal{X}:=\bigtimes_{i\in\mathcal{V}}\mathcal{X}_{i}$ and $\mathcal{U}:=\bigtimes_{i\in\mathcal{V}}\mathcal{U}_{i}$. Let $\bm{e}_{ij}:=\bm{x}_{j}-\bm{x}_{i}\in\mathbb{X}_{ij}$ represent the relative state vector for each $i,j\in\mathcal{V}$ where $\mathbb{X}_{ij}:=\mathbb{X}_{j}\ominus\mathbb{X}_{i}$ and let $S\in\mathbb{R}^{n_{p}\times n}$ be a selection matrix [^1] A selection matrix applies to select $m\leq n$ unique elements from a vector of dimensions $n$ such that $S\in\mathbb{R}^{m\times n}$ with $S[i,j]\in\{0,1\}$; $\sum_{j=1}^{n}S[i,j]=1,\,\forall i=1,\ldots m$ and $\sum_{i=1}^{m}S[i,j]=1,\,\forall j=1,\ldots n$. such that $\bm{p}_{i}=S\bm{x}_{i},\;\forall i\in\mathcal{V}$, represents the position vector of agent $i$ with dimension $n_{p}\leq n$.
II-A Polytopes and their properties
In this subsection, some fundamental properties of polytopes based on [32, Ch. 0-1] and [33, Sec. 3] are revised and adapted to the settings of this work for clarity of presentation.
@@ II-A Polytopes and their properties
Note that the inequality in the definition of $\mathcal{P}(A,\bm{c},\bm{z})$ should be interpreted row-wise. Moreover, the centre vector $\bm{c}$ is here introduced for the convenience of analysis, but the polytope definition $\mathcal{P}(A,\tilde{z})=\{\bm{x}\in\mathbb{R}^{n}|A\bm{x}\leq\tilde{\bm{z}}\}$ in [32, pp. 28] is equivalent to ours letting $\tilde{\bm{z}}=\bm{z}+A\bm{c}$.
@@
([32, pp. 4]) Given a finite set of points $V=\{\bm{v}_{k}\}_{k=1}^{|V|}\subset\mathbb{R}^{n}$ with $|V|\geq 1$, the convex hull of $V$ is defined as $conv(V):=\{\sum_{k=1}^{|V|}\lambda_{k}\bm{v}_{k}|\sum_{k=1}^{|V|}\lambda_{k}=1,\,\lambda_{k}\geq 0,\,\forall k=1,\ldots|V|\}$.
@@
Proposition 1 establishes three equivalent representations for a polytope $\mathcal{P}(A,\bm{c},\bm{z})$. Given $A$, $\bm{z}$ and $\bm{c}$, then a generator set $V$ can be obtained algorithmically (see for example [34]), while the functional representation $h(\bm{x})$ is obtained at no additional computational cost [32]. While for a given polytope the generator set $V$ is not unique, we define $\nu:\mathcal{P}(A,\bm{c},\bm{z})\mapsto V^{*}$ as the map between a polytope $\mathcal{P}(A,\bm{c},\bm{z})$ and the set of unique generators with minimal cardinality $V^{*}$ (the set of vertices for $\mathcal{P}$ [32, Prop. 2.2]). We next define when two polytopes are said to be similar and what properties can be derived from this similarity relation.
@@
Let the polytope $\mathcal{P}(\bm{A},0,\bm{z})$, the center vector $\bm{c}\in\mathbb{R}^{n}$ and scale $\alpha>0$ such that $\bm{\eta}=[\bm{c},\alpha]\in\mathbb{R}^{n+1}$. Moreover, let $V=\nu(\mathcal{P}(A,\bm{0},\bm{z}))$ be the generator set, such that $V=\{\bm{v}_{k}\}_{k=1}^{|V|}$, then
@@
In other words, the generators of a polytope $\mathcal{P}(A,\bm{c},\alpha\bm{z})$ are directly computed by translating and scaling the generators of the similar polytope $\mathcal{P}(A,\bm{0},\bm{z})$ as represented in Fig. 1. The next proposition extends the result of Prop. 2 to the Minkowski sum of similar polytopes.
@@
Let $\mathcal{P}(A,\bm{c}_{i},\alpha_{i}\bm{z}),\,\forall i\in\mathcal{I}$ be similar polytopes for some index set $\mathcal{I}$ and $\alpha_{i}\in\mathbb{R}_{+}$, then
@@
Proof:
Follows from [33, Thm. 3.2] and Prop. 2. ∎
Next, Prop. 4 provides a set of linear inequalities that can be employed to check the intersection and inclusion of two polytopes.
@@
Let the polytopes $\mathcal{P}(A_{1},\bm{c}_{1},\alpha_{1}\bm{z}_{1})$, $\mathcal{P}(A_{2},\bm{c}_{2},\alpha_{2}\bm{z}_{2})$ such that $A_{1}\in\mathbb{R}^{m_{1}\times n},A_{2}\in\mathbb{R}^{m_{2}\times n}$ and let $\bm{\eta}_{1}=[\bm{c}_{1},\alpha_{1}]$ and $\bm{\eta}_{2}=[\bm{c}_{2},\alpha_{2}]$. Furthermore, let $V_{1}=\nu(\mathcal{P}(A_{1},\bm{0},\bm{z}_{1})=\{\bm{v}_{k}\}_{k=1}^{|V_{1}|}$. Then, the inclusion relation $\mathcal{P}(A_{1},\bm{c}_{1},\alpha_{1}\bm{z}_{1})\subseteq\mathcal{P}(A_{2},\bm{c}_{2},\alpha_{2}\bm{z}_{2})$ holds if and only if
@@
Proof:
Given in Appendix A. ∎
The section is concluded with the following result, whose relevance is clarified in Sec. V-A.
@@
Thus Prop. 5 supports the intuition that given a polytope and a Minkowski sum of polytopes similar to the former, then the Minkoswki sum is included in the original polytope only if the sum of the scale factors $\alpha_{i}$ is less than or equal to 1. The relevance of this fact is clarified in Section V-A.
II-B Signal Temporal Logic
Signal Temporal Logic (STL) is a predicate logic applied to formally define spatio-temporal behaviours (tasks) for state signals $\bm{x}(t)\in\mathcal{X}$. Let $h:\mathcal{D}\rightarrow\mathbb{R}$, with $\mathcal{D}\subseteq\mathbb{X}$ be a predicate function and let the boolean-valued predicate
$\mu(h(\bm{x})):=\begin{cases}\top&\text{if}\;h(\bm{x})\geq 0\\ \bot&\text{if}\;h(\bm{x})<0.\end{cases}$
Then, the recursively defined STL grammar applies to construct complex specifications from a set of elementary predicates as